OpenStudy (anonymous):

What is the slope of the line that passes through the points (16, –1) and (–4, 10)? A.-20/11 B.-11/20 C.11/20 D.20/11

OpenStudy (solomonzelman):

Use the slope formula. \(\Large\color{blue}{ \bf \frac{y_1-y_2}{x_1-x_2} }\)

OpenStudy (anonymous):

I tried putting it in a calculator but it doesn't work :/

OpenStudy (anonymous):

do it with pencil and paper, it is much easier

OpenStudy (anonymous):

especially since all your answers are given in terms of fractions, not decimals

OpenStudy (anonymous):

I'll try..

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

ok what do i do with the 1 and 2?

OpenStudy (anonymous):

ok lets go slow the one and two are subscripts, meaning "first" and "second" you do not compute with them

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(\color{red}{16},\color{blue}{ –1}),(\color{red}{x_1},\color{blue}{y_1})\]

OpenStudy (anonymous):

\[(\color{red}{-4},\color{blue}{ 10}),(\color{red}{x_2},\color{blue}{y_2})\]

OpenStudy (anonymous):

What do i do with the y and x sorry I'm stupid lol

OpenStudy (anonymous):

lets go real slow no, not stupid, it is not like you are born knowing this in this case you can say \(y_2=10,y_1=-1\)

OpenStudy (anonymous):

and that makes \[y_2-y_1=10-(-1)=11\] that is the numerator of your fraction

OpenStudy (anonymous):

then \[x_2=-4,x_1=16\] and so \[x_2-x_1=-4-16=-20\] that is the denominator of your fraction

OpenStudy (anonymous):

so it's B?

OpenStudy (anonymous):

therefore the slope is \[\frac{y_2-y_1}{x_2-x_1}=\frac{10-(-1)}{-4-16}=\frac{11}{-20}\]

OpenStudy (anonymous):

B is a letter, the answer is \[-\frac{11}{20}\] want to try another one?

OpenStudy (anonymous):

No i meant i have B is -11/20 look at my question

OpenStudy (anonymous):

yes, that is right

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i have one more

OpenStudy (anonymous):

k lets try it

OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 8) and (12, 1)? A.-9/7 B.-7/9 C.7/9 D.9/7

OpenStudy (anonymous):

you got an idea this time? it is identical to the last one i mean the procedure is identical \[\frac{1-8}{12-3}\]

OpenStudy (anonymous):

-7/9

OpenStudy (anonymous):

yes hope it is clear how to do them now more clear anyways

OpenStudy (anonymous):

Yes they are thank you

OpenStudy (anonymous):

yw